The tens digit of a 2-digit number is twice the units digit. The sum of the digits is 12. Find the original number.
Answer by sudhanshu_kmr(1152)


let unit digit = x, tens digit =2x
2x + x = 12
=> x =12/3 =4
so, original number is 84.